ID |
Date |
Author |
Type |
Category |
Subject |
2223
|
Mon Nov 9 19:09:09 2009 |
Jenne | Update | PEM | Noise floor of the Ranger Seismometer |
To estimate the noise floor of the Ranger, Rana and I locked the mass on the seismometer, so there will be no (aka minimal) signal from the motion of the ground in the pickup coil. We should be seeing primarily the noise of the readout electronics. We also put the Ranger on top of one of the foam lids from the Seismometer Isolation Boxes to further isolate from ground motion (this didn't change the signal noticeably).
In this plot, Green is the locked-mass-on-foam noise floor, blue is the regular spectra, with the SR560 AC coupled, and the red is the regular seismic spectra with the SR560 DC coupled. There doesn't seem to be a noticeable difference between blue and red (the spectra were taken at different times of day, with the red taken at night, when we generally expect things to be quieter). I'm leaving the SR560 DC coupled. (Rana had found it earlier this afternoon GND coupled....not sure yet why).
Also, we're not sure that the green curve is true readout noise, vs. how much of it is specifically due to the fact that the mass is locked down. Especially around a few hundred Hz, the green curve is much higher than the other 2, and at a few tens of Hz there is some weird peak action. However, this will be okay as a first-run noise estimate for the Ranger's noise floor.
The question at hand is: Do we need to redo any of the Ranger's readout electronics (i.e. replace the SR560 with a Pomona Box circuit) to lower the noise floor, or is it okay as-is? |
Attachment 1: Ranger_noiseFloor_Spectra.png
|
|
2222
|
Mon Nov 9 19:04:23 2009 |
rob | Update | Computers | OMC FE hosed |
Quote: |
It won't start--it just sits at Waiting for EPICS BURT, even though the EPICS is running and BURTed.
[controls@c1omc c1omc]$ sudo ./omcfe.rtl
cpu clock 2388127
Initializing PCI Modules
3 PCI cards found
***************************************************************************
1 ADC cards found
ADC 0 is a GSC_16AI64SSA module
Channels = 64
Firmware Rev = 3
***************************************************************************
1 DAC cards found
DAC 0 is a GSC_16AO16 module
Channels = 16
Filters = None
Output Type = Differential
Firmware Rev = 1
***************************************************************************
0 DIO cards found
***************************************************************************
1 RFM cards found
RFM 160 is a VMIC_5565 module with Node ID 130
***************************************************************************
Initializing space for daqLib buffers
Initializing Network
Waiting for EPICS BURT
|
From looking at the recorded data, it looks like the c1omc started going funny on the afternoon of Nov 5th, perhaps as a side-effect of the Megatron hijinks last week.
It works when megatron is shutdown. |
2221
|
Mon Nov 9 18:32:38 2009 |
rob | Update | Computers | OMC FE hosed |
It won't start--it just sits at Waiting for EPICS BURT, even though the EPICS is running and BURTed.
[controls@c1omc c1omc]$ sudo ./omcfe.rtl
cpu clock 2388127
Initializing PCI Modules
3 PCI cards found
***************************************************************************
1 ADC cards found
ADC 0 is a GSC_16AI64SSA module
Channels = 64
Firmware Rev = 3
***************************************************************************
1 DAC cards found
DAC 0 is a GSC_16AO16 module
Channels = 16
Filters = None
Output Type = Differential
Firmware Rev = 1
***************************************************************************
0 DIO cards found
***************************************************************************
1 RFM cards found
RFM 160 is a VMIC_5565 module with Node ID 130
***************************************************************************
Initializing space for daqLib buffers
Initializing Network
Waiting for EPICS BURT
|
2220
|
Mon Nov 9 18:27:30 2009 |
Alberto | Frogs | Computers | OMC DCPD Interface Box Disconnected from the power Supply |
This afternoon I inadvertently disconnected one of the power cables coming from the power supply on the floor next to the OMC cabinet and going to the DCPD Interface Box.
Rob reconnected the cable as it was before. |
2219
|
Mon Nov 9 16:32:36 2009 |
Alberto | Frogs | Environment | Shot of the white board yesterday before erasing |
Yesterday Rana and I needed some room on the white board in the Control Room. We had to erase some of the stuff present on the board despite the bif warning "Do Not Erase".
This is how it looked like before erasing.

|
Attachment 1: DSC_0980.JPG
|
|
2218
|
Mon Nov 9 15:21:37 2009 |
Jenne | Update | General | Drill Press is down for the count |
Quote: |
The on/off switch for the drill press is broken. Replacement parts should be here tomorrow.
|
Drill press is all better now. A spare switch is in the top drawer with the drill bits. |
2217
|
Mon Nov 9 15:11:02 2009 |
Alberto | Update | LSC | Everything put back as it was |
Quote: |
I disconnected the setup for the arm cavity TF measurement. I opened the scitation switch on the ISS medm screen. I reconnected the OMC ISS EXC cable to the breakout box on the floor.
The photodiode on the Y end is stilll connected.
Also the RFAm (whish is not disbaled anymore) still has a 50% beam splitter before it.
I'm also running the Align Full IFO script.
|
I removed the beam splitter and the PDA 255.
the beam path to the RFAM photodiode is clear again. |
2216
|
Mon Nov 9 15:08:29 2009 |
Koji | Omnistructure | Environment | Tidying up BNC cables rack around the lab |
Quote: |
This would be a good trial once you put the label "BNC only" on the wall.
Quote: |
We have thousands of miles of BNC cables in the lab but we still don't find one when we need it. I decided to solve the problem.
This morning I tried to tidy up the several cable rack the we have in the lab.
i tried to dedicate each rack to a speecific type of cables: special cables, hand made cables, BNCs, LEMOs, etc.
In particular I tryed to concentrate BNC cable of several lengths on the rack near by the ITMX chamber.
People are invited to preserve the organization.
|
|
Done! Check it out. |
2215
|
Mon Nov 9 14:59:34 2009 |
josephb, alex | Update | Computers | The saga of Megatron continues |
Apparently the random file system failure on megatron was unrelated to the RFM card (or at least unrelated to the physical card itself, its possible I did something while installing it, however unlikely).
We installed a new hard drive, with a duplicate copy of RTL and assorted code stolen from another computer. We still need to get the host name and a variety of little details straightened out, but it boots and can talk to the internet. For the moment though, megatron thinks its name is scipe11.
You still use ssh megatron.martian to log in though.
We installed the RFM card again, and saw the exact same error as before. "NMI EVENT!" and "System halted due to fatal NMI".
Alex has hypothesized that the interface card the actual RFM card plugs into, and which provides the PCI-X connection might be the wrong type, so he has gone back to Wilson house to look for a new interface card. If that doesn't work out, we'll need to acquire a new RFM card at some point
After removing the RFM card, megatron booted up fine, and had no file system errors. So the previous failure was in fact coincidence.
|
2214
|
Mon Nov 9 14:53:47 2009 |
Alberto | Omnistructure | Environment | Tidying up BNC cables rack around the lab |
This would be a good trial once you put the label "BNC only" on the wall.
Quote: |
We have thousands of miles of BNC cables in the lab but we still don't find one when we need it. I decided to solve the problem.
This morning I tried to tidy up the several cable rack the we have in the lab.
i tried to dedicate each rack to a speecific type of cables: special cables, hand made cables, BNCs, LEMOs, etc.
In particular I tryed to concentrate BNC cable of several lengths on the rack near by the ITMX chamber.
People are invited to preserve the organization.
|
|
2213
|
Mon Nov 9 13:26:19 2009 |
Alberto | Omnistructure | Environment | Tidying up BNC cables rack around the lab |
We have thousands of miles of BNC cables in the lab but we still don't find one when we need it. I decided to solve the problem.
This morning I tried to tidy up the several cable rack the we have in the lab.
i tried to dedicate each rack to a speecific type of cables: special cables, hand made cables, BNCs, LEMOs, etc.
In particular I tryed to concentrate BNC cable of several lengths on the rack near by the ITMX chamber.
People are invited to preserve the organization.
|
Attachment 1: DSC_0987.JPG
|
|
2212
|
Mon Nov 9 13:22:08 2009 |
josephb,alex | Update | Computers | Megatron update |
Alex and I took a look at megatron this morning, and it was in the same state I left it on Friday, with file system errors. We were able to copy the advLIGO directory Peter had been working in to Linux1, so it should be simple to restore the code. We then tried just running fsck, and overwritting bad sectors, but after about 5 minutes it was clear it could potentially take a long time (5-10 seconds per unreadable block, with an unknown number of blocks, possibly tens or millions). The decision was made to simply replace the hard drive.
Alex is of the opinion that the hard drive failure was a coincidence. Or rather, he can't see how the RFM card could have caused this kind of failure.
Alex went to Bridge to grab a usb to sata adapter for a new hard drive, and was going to copy a duplicate install of the OS onto it, and we'll try replacing the current hard drive with it. |
2211
|
Mon Nov 9 13:17:07 2009 |
Alberto | Configuration | ABSL | NPRO on |
I turned the auxialiary NPRO for the AbsL Experiment on. Its shutter stays closed. |
2210
|
Mon Nov 9 12:09:10 2009 |
Alberto | Omnistructure | Environment | BNC Cable Laid Down from South End to East VErtex |
I laid down the floor a BNC cable from the Y End table to the BNC Chamber. The cable runs next to the east wall.
I'm leaving the cable because it can turn useful in the future.
I'm tying the end of the cable to a big threaded steel rod on the side of the BS chamber.
I've also labeled as TRY |
Attachment 1: DSC_0986.JPG
|
|
2209
|
Mon Nov 9 11:14:57 2009 |
Alberto | Update | ABSL | Started working on the PSL table |
I'm working on the PSL table to set up the PLL setup for the AbsL experiment. |
2208
|
Mon Nov 9 11:11:19 2009 |
Alberto | Configuration | LSC | MC2 Watchdogs tripped |
The MC2 watchdogs tripped. I just restored them.
I also had to relock the MZ and then the Mode Cleaner. |
2207
|
Mon Nov 9 08:43:16 2009 |
steve | Update | SUS | MC2 damping restored |
MC2 damping restored, MZ locked and the arms are flashing now. |
2206
|
Mon Nov 9 01:52:56 2009 |
rana | Update | PEM | coherence v. time for 2 accelerometers |
I used the coh_carpet.m function from the mDV to calculate this plot:
coh_carpet('C1:PEM-ACC_MC1_X','C1:PEM-ACC_MC2_X',gps('now - 3 days'),3600*12,4,10,64)
It shows the coherence v. time of two of our X-direction accelerometers starting around 1AM on Friday and going for 12 hours.
I'm not sure what it means exactly, but it looks like the coherence is relatively steady as a function of time. I will need more RAM than Rosalba or a smarter code to calculate longer time stretches. |
Attachment 1: coh.png
|
|
2205
|
Sun Nov 8 22:50:29 2009 |
Alberto | Update | ASC | IFO Alignment |
Tonight I aligned the IFO by running the scripts one by one.
SRC was far off and I had to align SRM by hand before the script could work. SPOB is still low when DRM is aligned.
I'm restoring the full IFO now that I'm taking off. |
2204
|
Sun Nov 8 14:18:25 2009 |
Alberto | Update | SUS | ETMY Watchdogs tripped |
This afternoon I re-enabled the ETMY coils after I found that the watchdogs for the mirror had tripped last night at 2:06am. |
2203
|
Sat Nov 7 23:50:45 2009 |
Haixing | Update | General | Open-loop transfer function of the magnetic levitation system |
I measured the open-loop transfer function of the magnetic levitation system.
The schematic block diagram for this measurement is the following:

I injected a signal at a level of 20mV between two preamplifiers, and the corresponding open-loop
transfer function is given by B/A. I took a picture of the resulting measurement, because
I encountered some difficulties to save the data to the computer via the wireless network.
The bode plots for the transfer function shown on the screen is the following:

I am puzzled with the zero near 10 Hz. I think it should come from the mechanical response function, because there is no zero in the transfer functions
of the preamplifer and the coil itself. I am not sure at the moment.
The corresponding configuration of the levitated magnet is

|
2202
|
Fri Nov 6 23:02:44 2009 |
Haixing | Update | General | SR785 Spectrum Analyzer |
I am using SR785 Spectrum Analyzer now and also tomorrow.
I will put it back on Sunday. If anyone needs it during the weekend,
please just take it and I can reset it by myself later. Thanks. |
2201
|
Fri Nov 6 20:10:15 2009 |
Jenne | Update | elog | elog acting up |
Quote: |
elog was acting up again (not running), so I restarted it.
|
And again. This makes 4 times since lunchtime yesterday....something bad is up. |
2200
|
Fri Nov 6 19:29:24 2009 |
Jenne | Update | elog | elog acting up |
elog was acting up again (not running), so I restarted it. |
2199
|
Fri Nov 6 19:25:31 2009 |
Sanjit, Jenne, Joe | Update | Adaptive Filtering | More work on saving coeffs on the OAF screen |
Quote: |
I made some changes in the code (all commented in the installed and SVN version) to print the filter coefficients. I got crazy output. Sometimes memory bugs lead to such crazy behavior. So far I could not find any bugs, but will have to spend more time on it
|
Something strange was going on in the OAF code, printf would print a double precision number in %f format but not in %lf or %e format!
Since we know this problem now, we can move forward, but it will be important to know why printf was restricted and if there are other such constraints which we should remember while making changes in the codes.
|
2198
|
Fri Nov 6 18:52:09 2009 |
pete | Update | Computers | RCG ETMY plan |
Koji, Joe, and I are planning to try controlling the ETMY, on Monday or Tuesday. Our plan is to try to do this with megatron out of the RFM loop. The RCG system includes pos, pit, yaw, side, and oplevs. I will use matrix elements as currently assigned (i.e. not the ideal case of +1 and -1 everywhere). I will also match channel names to the old channels. We could put buttons on the medm screen to control the analog DW by hand.
This assumes we can get megatron happy again, after the unhappy RFM card test today. See Joe's elog immediately before this one.
We first plan to test that the ETMY watchdog can disable the RCG frontend, by using a pos step (before connecting to the suspension). Hopefully we can make it work without the RFM. Otherwise I think we'll have to wait for a working RFM card.
We plan to disable the other optics. We will disable ETMY, take down the ETMY frontend, switch the cables, and start up the new RCG system. If output looks reasonable we will enable the ETMY via the watchdog. Then I suppose we can put in some small steps via the RCG controller and see if it damps.
Afterwards, we plan to switch everything back. |
2197
|
Fri Nov 6 18:13:34 2009 |
josephb | Update | Computers | Megatron woes |
I have removed the RFM card from Megatron and left it (along with all the other cables and electronics) on the trolly in front of the 1Y9 rack.
Megatron proceeded to boot normally up until it started loading Centos 5. During the linux boot process it checks the file systems. At this point we have an error:
/dev/VolGroup00/LogVol00 contains a file system with errors, check forced
Error reading block 28901403 (Attempt to read block from filesystem resulted short read) while doing inode scan.
/dev/VolGroup00/LogVol00 Unexpected Inconsistency; RUN fsck MANUALLY
So I ran fsck manually, to see if I get some more information. fsck reports back it can't read block 28901403 (due to a short read), and asks if you want to ignore(y)?. I ignore (by hitting space), and unfortunately touch it an additional time. The next question it asks is force rewrite(y)? So I apparently forced a rewrite of that block. On further ignores (but no forced rewrites) I continue seeing short read errors at 28901404, *40, *41,*71, *512, *513, etc. So not totally continugous. Each iteration takes about 5-10 seconds. At this point I reboot, but the same problem happens again, although it starts 28901404 instead of 28901403. So apparently the force re-write fixed something, but I don't know if this is the best way of going about this. I just wondering if there's any other tricks I can try before I just start rewriting random blocks on the hard drive. I also don't know how widespread this problem is and how long it might take to complete (if its a large swath of the hard drive and its take 10 seconds for each block that wrong, it might take a while).
So for the moment, megatron is not functional. Hopefully I can get some advice from Alex on Monday (or from anyone else who wants to chime in). It may wind up being easiest to just wipe the drive and re-install real time linux, but I'm no expert at that.
|
2196
|
Fri Nov 6 18:02:22 2009 |
josephb | Update | Computers | Elog restarted |
While I was writing up an elog entry, the elog died again, and I restarted it. Not sure what caused it to die since no one was uploading to it at the time. |
2195
|
Fri Nov 6 17:04:01 2009 |
josephb | Configuration | Computers | RFM and Megatron |
I took the RFM 5565 card dropped off by Jay and installed it into megatron. It is not very secure, as it was too tall for the slot and could not be locked down. I did not connect the RFM fibers at this point, so just the card is plugged in.
Unfortunately, on power up, and immediately after the splash screen I get "NMI EVENT!" and "System halted due to fatal NMI".
The status light on the RFM light remains a steady red as well. There is a distinct possibility the card is broken in some way.
The card is a VMIPMC-5565 (which is the same as the card used by the ETMY front end machine). We should get Alex to come in and look at it on Monday, but we may need to get a replacement. |
2194
|
Fri Nov 6 16:27:15 2009 |
Jenne | Update | PEM | Ranger moved |
The Ranger seismometer has been moved to ~the middle of the Mode Cleaner tube, and it's orientation has been changed to horizontal (using all of the locking/mass centering procedures). This is similar in orientation to the way things were back in the day when Rana and Matt had the OAF running nicely. |
2193
|
Fri Nov 6 12:56:30 2009 |
Haixing | Update | SUS | Magnet has been levitated |
In this experiment, we used a feedback control to create a stable trap for a NdFeB permanent magnet. The block diagram is the following:

The displacement of the magnet is sensed by the Hall-effect sensor, whose output voltage is proportional to the magnetic flux produced
by the permanent magnet. It has a flat response within the frequencies we are interested in. It is driven by a 5 V power supplier and its
output has a DC voltagle of 2.5 V. We subtracted the DC voltage and used the resulting signal as the error signal. This was
simply achieved by using two channels "A" and "B". The output is "A-B" with a gain equal to one. We then put the error
signal into another SR560 as a low-pass filter with a gain of 100 above 30 Hz. We used the "DC" coupling modes in both
preamplifers. The output is then used to drive a coil. The coil has a dimension of 1.5 inch in diameter and 2 inch in length.
The inductance of the coil is around 0.5 H and the resistance is 4.7 Om. Therefore, it has a corner frequency aournd 10/2pi Hz.
The coil has a iron core inside to enhance the DC force to the permanent magnet. The low-frequency 1/f response of the magnet is produced
by the eddy current damping of the aluminum plane that is below the magnet. This 1/f response is essential for a stable configuration. In the
next stage, we will remove the aluminum plane, and instead we will use a filter to create similar transfer function. At high-frequencies, it behaves as
a free-mass and has a 1/f^2 response. Finally, the magnet is stably levitated.
|
Attachment 1: DSC_0964.JPG
|
|
2192
|
Fri Nov 6 10:35:56 2009 |
josephb | Update | Computers | RFM reboot fest and re-enabled ITMY coil drivers |
As noted by Steve, the RFM network was down this morning. I noticed that c1susvme1 sync counter was pegged at 16384, so I decided to start with reboots in that viscinity.
After power cycling crates containing c1sosvme, c1susvme1, and c1susvme2 (since the reset buttons didn't work) only c1sosvme and c1susvme2 came back normally. I hooked up a monitor and keyboard to c1susvme1, but saw nothing. I power cycled the c1susvme crate again, and this time I watched it boot properly. I'm not sure why it failed the first time.
The RFM network is now operating normally. I have re-enabled the watchdogs again after having turned them off for the reboots. Steve and I also re-enabled the ITMY coil drivers when I noticed them not damping once the watch dogs were re-enabled. The manual switches had been set to disabled, so we re-enabled them. |
2191
|
Fri Nov 6 09:17:53 2009 |
steve | Omnistructure | PEM | PSL HEPAs turned on |
The PSL enclosure HEPAs turned on at 30% |
2190
|
Fri Nov 6 07:55:59 2009 |
steve | Update | Computers | RFMnetwork is down |
The RFMnetwork is down. MC2 sus damping restored. |
2189
|
Fri Nov 6 00:40:29 2009 |
Alberto | Update | LSC | Everything put back as it was |
I disconnected the setup for the arm cavity TF measurement. I opened the scitation switch on the ISS medm screen. I reconnected the OMC ISS EXC cable to the breakout box on the floor.
The photodiode on the Y end is stilll connected.
Also the RFAm (whish is not disbaled anymore) still has a 50% beam splitter before it.
I'm also running the Align Full IFO script. |
2188
|
Fri Nov 6 00:24:06 2009 |
Alberto | Update | LSC | Y Arm Cavity Transfer Function |
Quote: |
As for the X Arm, this the transfer function I measured for the Y arm cavity.

This time I'm using a different photodiode than the PDA255 on the Y end table.
The diode I'm using is the PDA520 from where TRY comes from.
I'm going to repeat the measurement with PDA255.
|
Measurement repeated with the PDA255 PD at the end but not big changes

|
2187
|
Fri Nov 6 00:23:34 2009 |
Alberto | Configuration | Computers | Elog just rebooted |
The elog just crashed and I rebooted it |
2186
|
Thu Nov 5 23:09:34 2009 |
Alberto | Update | LSC | Y Arm Cavity Transfer Function |
As for the X Arm, this the transfer function I measured for the Y arm cavity.

This time I'm using a different photodiode than the PDA255 on the Y end table.
The diode I'm using is the PDA520 from where TRY comes from.
I'm going to repeat the measurement with PDA255. |
2185
|
Thu Nov 5 22:30:09 2009 |
Alberto | Update | LSC | X Arm Cavity Transfer Function |
It seems that just repeating the measurement was enough to get a good transfer function of the x arm cavity. Here's what I got.

I'm going to fit the data on matlab, but at first sight, the pole seems to be at about 1.7KHz (that is where the phase is 45deg): as expected.
Probably it was useful to realign the beam on the Transmission PD. (btw, I'm using the PDA255 that was still on the X end table since the AbsL experiemtn that measured the arm length) |
2184
|
Thu Nov 5 19:25:11 2009 |
Alberto | Update | LSC | X Arm Cavity transfer Function |
Quote: |
I would have guessed that you have to calibrate the detectors relative to each other before trying this. Its also going to be tricky if you use 2 different kinds of ADC for this (c.f. today's delay discussion in the group meeting).
I think Osamu used to look at fast transmission signals by making sure the PD at the end had a 50 Ohm output impedance and just drive the 40m long cable and terminate the receiving end with 50 Ohms. Then both PDs go into the SR785.
|
On Rana's suggestion I measured the trasfer function between the two photodiodes PDA255 that I'm using.
I took the one that I had on the end table and put it on the PSL table. I split the MC transmitted beam with a 50% beam splitter and sent the beams on the two diodes. (Rana's idea of picking off the beam and interposing the PDs before the ISS PDs was not doable: ISS PDs would be too close and there would be no room to install the PDA255 before them). See picture with the final setup.

The transfer function also includes the 40m long cable that I was using for the Arm Cavity measurement.
Here's what I got. It looks rather flat. Yesterday the calibration was probably not the problem in that measurement.

I'm now going to install the PD back on the end table and measure the TFs between the excitation and several points of the loop.
(Trivia. At first, the PDs were saturating so Koji attached attenuation filters on to them. Suddenly the measurement got much nicer) |
2183
|
Thu Nov 5 16:41:14 2009 |
josephb | Configuration | Computers | Megatron's personal network |
In investigating why megatron wouldn't talk to the network, I re-discovered the fact that it had been placed on its own private network to avoid conflicts with the 40m's test point manager. So I moved the linksys router (model WRT310N V2) down to 1Y9, plugged megatron into a normal network port, and connected its internet port to the rest of the gigabit network.
Unfortunately, megatron still didn't see the rest of the network, and vice-versa. I brought out my laptop and started looking at the settings. It had been configured with the DMZ zone on for 192.168.1.2, which was Megatron's IP, so communications should flow through the router. Turns out it needs the dhcp server on the gateway router (131.215.113.2) to be on for everyone to talk to each other. However, this may not be the best practice. It'd probably be better to set the router IP to be fixed, and turn off the dhcp server on the gateway. I'll look into doing this tomorrow.
Also during this I found the DNS server running on linux1 had its IP to name and name to IP files in disagreement on what the IP of megatron should be. The IP to name claimed 131.215.113.95 while the name to IP claimed 131.215.113.178. I set it so both said 131.215.113.178. (These are in /var/named/chroot/var/ directory on linux1, the files are 113.215.131.in-addr.arpa.zone and martian.zone - I modified the 113.215.131.in-addr.arpa.zone file). This is the dhcp served IP address from the gateway, and in principle could change or be given to another machine while the dhcp server is on. |
2182
|
Thu Nov 5 16:30:56 2009 |
pete | Update | Computers | moving megatron |
Joe and I moved megatron and its associated IO chassis from 1Y3 to 1Y9, in preparations for RCG tests at ETMY. |
2181
|
Thu Nov 5 16:24:59 2009 |
Koji | Update | CDS | ETMY CDS test stuff |
Joe, Peter, Jay, Koji, Rana
We put the new CDS stuff at Y end 1Y9 rack.
Items
- megatron
- wireless router
- IO chasis (black)
- Extention cable (between megatron & IO chasis)
- 1 ADC card
- 1 DAC card
- 1 BIO card
- The adapter box for ADC
- The adapter box for DAC
- The adapter box for BIO
- 2x IDC-DB37 cable for the ADC box - AA chasis
- 1x IDC cable for the DAC box - Pentek
- 1x DB cable for the BIO box
- 1x +/-15V cable for the BIO box
|
2180
|
Thu Nov 5 16:24:40 2009 |
Jenne | Update | General | Drill Press is down for the count |
The on/off switch for the drill press is broken. Replacement parts should be here tomorrow. |
2179
|
Thu Nov 5 12:34:26 2009 |
kiwamu | Update | Computers | elog rebooted |
I found elog got crashed. I rebooted the elog daemon just 10minutes before. |
2178
|
Thu Nov 5 05:07:22 2009 |
rana | Update | LSC | X Arm Cavity transfer Function |
I would have guessed that you have to calibrate the detectors relative to each other before trying this. Its also going to be tricky if you use 2 different kinds of ADC for this (c.f. today's delay discussion in the group meeting).
I think Osamu used to look at fast transmission signals by making sure the PD at the end had a 50 Ohm output impedance and just drive the 40m long cable and terminate the receiving end with 50 Ohms. Then both PDs go into the SR785.
|
2177
|
Wed Nov 4 23:17:51 2009 |
Alberto | Update | LSC | X Arm Cavity transfer Function |
I measured the transfer function between MC_TRANS and TRX and I'm attaching the result.

That looks quite strange. Something's wrong. I'll repeat it tomorrow.
for the night I'm putting everything back. I'm also reconnecting the OMC_ISS_EXC and opening again the test switch on the ISS screen.
The RFAM monitor remains disable |
2176
|
Wed Nov 4 21:46:18 2009 |
Alberto | Update | LSC | Arm Cavity Finesse Measurement |
Quote: |
Quote: |
I'm going to work on the X arm to measure the arm cavity finesse.
The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.
I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done.
|
Since I need to measure the transfer function between TRX and MC_TRANS_DC I picked off the beam going to RFAM PD to send it to a PDA255 photodiode (cannibalized from the AbsL's PLL) which I installed on the PSL table.
I centerd the beam on the PD and I was able to see the injected signal.
I think I'm ready to measure the transfer function.
Except for the RFAM PD everything is as before.
I'm gonna go grab dinner and I should be back to keep working on that in about one hour.
|
Back from dinner. Taking measurements. |
2175
|
Wed Nov 4 18:35:19 2009 |
Alberto | Update | LSC | Arm Cavity Finesse Measurement |
Quote: |
I'm going to work on the X arm to measure the arm cavity finesse.
The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.
I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done.
|
Since I need to measure the transfer function between TRX and MC_TRANS_DC I picked off the beam going to RFAM PD to send it to a PDA255 photodiode (cannibalized from the AbsL's PLL) which I installed on the PSL table.
I centerd the beam on the PD and I was able to see the injected signal.
I think I'm ready to measure the transfer function.
Except for the RFAM PD everything is as before.
I'm gonna go grab dinner and I should be back to keep working on that in about one hour. |
2174
|
Wed Nov 4 16:49:32 2009 |
Alberto | Update | LSC | Arm Cavity Finesse Measurement |
I'm going to work on the X arm to measure the arm cavity finesse.
The idea is to measure the cavity transfer function to estimate the frequency of its cavity pole. That should be a more accurate measurement than that based on the cavity decay time.
I'm starting now and I'm going to inject a swept sine excitation on the OMC_ISS_EXC input cable laying on the floor nearby the AP table (see pic).

In orderf to do that I disconnected the cable from the OMC breakout box laying on the floor. I'm going to plug the cable back in as soon as I'm done. |